OpenStudy (anonymous):

the square root of negative 689

OpenStudy (anonymous):

26.248809

OpenStudy (anonymous):

\[\sqrt{-689}\] \[\sqrt{689}i\]

OpenStudy (anonymous):

there is no real square root of a negative number. you can write \[\sqrt{689}i\] if you want